OpenStudy (anonymous):

Which equation is a quadratic equation? y + 3 = -2x + 5 - 8x y + 2x = -2x(4 - 3x) + 4 6x3 + y = (2x - 2)(3x + 7) 9x - y = (3x2 + 1)(5x - 1) + 4x

OpenStudy (anonymous):

can u write the equations separately ?

OpenStudy (anonymous):

excuse me @UserA113 post ur question separately please

OpenStudy (anonymous):

i dont know how to post a picture tho :/

OpenStudy (anonymous):

dont post the pic in the ques its impossible so first post ur ques then post the pic in the reply column just like u did in this ques

OpenStudy (anonymous):

Which equation is a quadratic equation? y + 3 = -2x + 5 - 8x y + 2x = -2x(4 - 3x) + 4 6x3 + y = (2x - 2)(3x + 7) 9x - y = (3x2 + 1)(5x - 1) + 4x

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

@UserA113

OpenStudy (anonymous):

uhm... Can you help with mine?

OpenStudy (anonymous):

ya sure let me think first :)

OpenStudy (anonymous):

oh thank you!

OpenStudy (anonymous):

in the 3rd equa is that \[6^3\]or 6x3

OpenStudy (anonymous):

6x^3

OpenStudy (anonymous):

so i dont think any of these is a quadratic equation because quadratic equation is something \[x^2+11x+28=0\]

OpenStudy (anonymous):

those are the answers though..what should I do?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

oh Im screwed :(

OpenStudy (aum):

\[\Large y + 3 = -2x + 5 - 8x \\ \Large y + 2x = -2x(4 - 3x) + 4 \\ \Large 6x^3 + y = (2x - 2)(3x + 7) \\ \Large 9x - y = (3x^2 + 1)(5x - 1) + 4x \\ \]The highest power of x in the first one is just 1. (not a quadratic) The highest power of x in the third one is 3 due to x^3 on the LHS. (not a quadratic) If you mentally multiply it out, the highest power in the fourth one is 3 which comes from multiplying 3x^2 and 5x which is 15x^3 (not a quadratic). But in the second case, the highest power is 2 which comes from multiplying (-2x) and (-3x) which is 6x^2. This is a quadratic equation.

OpenStudy (anonymous):

oh That makes sense! Thank you! Can you help me with another one?

OpenStudy (aum):

I can try one more.

OpenStudy (anonymous):

Identify the vertex for the graph of y = -2x2 - 12x + 5. (3, -13) (-3, -13) (-3, 23) (3, 23)

OpenStudy (aum):

In ax^2 + bx + c, the x-coordinate of the vertex is -b / (2a). So what will be the x-coordinate of the vertex in -2x^2 - 12x + 5 ?

OpenStudy (anonymous):

2?

OpenStudy (aum):

-2x^2 - 12x + 5 a = -2, b = -12, c = 5 what is -b / (2a)?

OpenStudy (anonymous):

@aum can u help me with my ques

OpenStudy (anonymous):

12? so the answer is 6a?

OpenStudy (aum):

I can try after I finish this one.

OpenStudy (anonymous):

oh...I don't mean to be rude but after sheelo can you keep helping me? Because you help a lot

OpenStudy (aum):

-2x^2 - 12x + 5 a = -2, b = -12, c = 5 -b / (2a) = -(-12) / (2 * -2) = 12 / -4 = -3. Put x = -3 in -2x^2 - 12x + 5 to find y.

OpenStudy (anonymous):

ya its ok @aum can help me after he/she finishes ur ques

OpenStudy (anonymous):

11x^2?

OpenStudy (aum):

@sheelo_mughal : Looks like others have already answered the questions you have posted si far in that link. If you have more post them there and I will take a look. @lehmonyo : I will have to log off shortly after answering her question.

OpenStudy (anonymous):

oh alright :(

OpenStudy (aum):

Put x = -3 in -2x^2 - 12x + 5 to find y. -2(-3)^2 - 12(-3) + 5 = ?

OpenStudy (anonymous):

I thought it was 11x squared

OpenStudy (aum):

Where did you get the 11x^2 from?

OpenStudy (anonymous):

from the equation

OpenStudy (aum):

In the problem the given equation is: y = -2x^2 - 12x + 5. We found the x value from -b/2a which is -(-12) / (2*(-2)) = 12/-4 = -3 We need to find y. Put x = -3 in y = -2x^2 - 12x + 5. y = -2(-3)^2 - 12(-3) + 5 y = -18 + 36 + 5 = 23 So the vertex is at (-3, 23)

OpenStudy (anonymous):

ohhh alright thank you. Do you have to leave now or..?